# Trajectory-to-Action Pipeline (TAP): Automated Scenario Description Extraction for Autonomous Vehicle Behavior Comparison

### v1

@@ 1. Introduction
Autonomous vehicles (AVs) are nearing widespread deployment, but ensuring their safety remains a critical challenge. Public AV testing has revealed critical failures, sometimes leading to tragic outcomes (mccarthy2022autonomous, ; shepardson_2024, ; brown2023halting, ; bote_2023, ; song_2021, ; templeton_2023, ; o2016we, ; rodriguez_2023, ). A methodical approach for measuring and improving safety for AVs is necessary to reduce these outcomes. A central question in AV development and testing is: How do we know which autonomous vehicle (AV) is safer? This fundamental question underlines much of the debate surrounding AV deployment and safety. Conventional metrics such as miles driven or disengagement counts provide a limited and often inadequate view of AV performance. Lower disengagements, for instance, do not necessarily equate to safer or more reliable driving behavior, as it fails to capture the full range of complex interactions AVs encounter in real-world traffic situations. To objectively compare the safety of two AVs, we need a framework that goes beyond these basic metrics.
@@
Comparing the safety of two AVs is challenging because they often operate under different conditions, different cities etc. Direct safety comparisons are unreliable without a common basis. Rather, a more viable approach is to analyze AV behavior when exposed to similar traffic scenarios. By evaluating AV performance in high-stakes situations, we gain deeper insight into their safety and robustness. However, implementing and scaling this approach requires automatically describing, interpreting, and comparing traffic scenarios across large datasets. This challenge motivates our work: How can we generate standardized, interpretable descriptions of traffic scenarios to identify both similar and rare situations for AV comparison?
@@
Scenario Description Languages (SDLs) provide structured, interpretable descriptions of traffic scenarios by capturing both static (road geometry, traffic signals, etc.) and dynamic (vehicle maneuvers) elements. SDLs can facilitate critical AV testing and safety tasks such as large-scale scenario discovery, real-to-sim transfers, and programmatic scenario generation. Additionally, by encoding traffic scenarios into SDL embeddings, we can identify similar scenarios allowing us to observe AVs in a similar traffic situation and, hence, reasoning about their relative safety. Frameworks such as Scenic (Scenic, ) and OpenScenario (OpenScenario, ) offer SDLs for these purposes.
@@
However, despite their utility, significant limitations of constructing SDLs remain: Due to the lack of ground-truth annotations, the extraction of semantic SDL labels from AV data is currently manual, requiring intensive oversight. Manual labeling of such data is impractical due to the vast scale of AV trajectory datasets. Since AV data sets do not provide labeled actions, using supervised learning techniques to extract SDLs is infeasible. This bottleneck limits the scalability of SDLs for comprehensive scenario-based testing across diverse AV datasets. While advances in computer vision have made static element extraction more feasible (segment_anything, ; SemSeg_Transformer, ), extracting dynamic vehicle behaviors is far more complex. Accurately capturing spatio-temporal actions requires more than video-based similarity, as it involves understanding nuanced trajectory patterns.
@@
To overcome these limitations, we introduce the Trajectory-to-Action Pipeline (TAP), a method that automates the extraction of interpretable vehicle behavior SDL descriptions from unlabeled AV trajectory data. TAP uses cross-entropy optimization to automatically derive SDL labels based on the dataset’s distribution. This automated approach addresses the pressing need for scalable, interpretable scenario descriptions, enabling TAP to compare vehicle behaviors and identify unique vehicle behaviors without the need for manual input. The contributions of this paper are as follows:
  1. (1)
@@ 2. Related Work
Scenario Description Languages Scenario Description Languages (SDLs) serve as high-level embeddings to describe traffic scenarios, capturing key elements like vehicles, pedestrians, and traffic lights. Existing SDLs such as Fortellix’s MSDL (M-SDL, ), Scenic (Scenic, ), and OpenScenario (OpenScenario, ) focus on comprehensive scene descriptions of environment, road geometry etc. However, our work diverges by concentrating specifically on describing autonomous vehicle motion to facilitate behavior comparison. Previous methods, such as Scenario2Vector (aron_date24, ; aron_iccps21, ), have explored using SDLs for scenario similarity using video and natural language data. In contrast, our approach differs by introducing an automated pipeline for extracting vehicle motion - focused SDL labels, enhancing scalability and precision. Alternative approaches include graph-based representations like Roadscene2vec and GraphAD (Roadscene2vec, ; GraphAD, ; PreGSU, ), which model scenes using nodes for agents and edges for their interactions. While effective for capturing vehicle interactions, these methods struggle to accurately represent vehicle motion dynamics. Recently, natural language descriptions (scene_understanding_captions, ; explainable_cars, ) have been used to capture complex scene interactions. Although expressive, this approach requires extensive training data and includes extraneous linguistic elements, making it less efficient for scenario comparison. In contrast, our method achieves a high information density while maintaining interpretability, focusing on the essential motion characteristics relevant to vehicle behavior.
@@
Scene Understanding Scene understanding is a vast field within computer vision, of which traffic scene understanding is a small portion. In (nlp_video_retrieval1, ), researchers use dashcam video for scenario retrieval. Other scene understanding work (SpatialCNN_SceneUnderstanding, ; RoadSceneUnderstanding, ; Vid2Bev, ) focuses on modeling the relationship between road agents and their environment features like lanes, etc. For instance, Vid2Bev (Vid2Bev, ) converts camera footage into a bird’s-eye view, enabling AVs to create a vectorized world for planning and control. These algorithms are important for AVs to create a vectorized view of the world in which their planning and control modules operate. Since AV planning and control modules already operate in this vectorized space, we leverage the existing data available to the vehicle. This allows us to analyze vehicle behavior as perceived by the vehicle itself, rather than relying on camera-based perception. Thus, our focus is on analyzing vectorized vehicle data, bypassing the need to re-implement perception algorithms.
@@
Trajectory Similarity In the vectorized world used by AVs, vehicle behavior is represented by trajectories of positional history. Traditional methods for trajectory similarity, such as Average Displacement Error (ADE) (ADE, ) and Dynamic Time Warping (DTW) (DTW, ), compare distances between corresponding points. However, these methods struggle with varying trajectory lengths, requiring resampling, which can strip away crucial information. Deep learning methods (dnn_traj_sim, ) are more robust to length differences but remain sensitive to early deviations, leading to accumulated errors. In contrast, our Trajectory-to-Action Pipeline (TAP) uses Scenario Description Language (SDL) labels to capture interpretable action-level behaviors, preserving essential motion characteristics and ensuring robustness to trajectory variations.
3. Scenario Description Language
@@
3.1. Overview of SDL Structure
Figure 1. Actions in our SDL, organized along a spectrum of complexity. At the abstract level, there are three lateral and three longitudinal labels, which can be further subdivided into finer-grained labels for more detailed analysis.
SDLs provide a systematic way to describe driving behaviors. Based on the SDL framework outlined by (aron_date24, ), we define an SDL that characterizes autonomous vehicle behaviors through a defined set of temporal actions. Figure 1 illustrates the set of vehicle behaviors employed in our SDL framework. We partition the driving behavior of the autonomous vehicle in terms of lateral and longitudinal actions, where Lateral actions capture directional movements, such as turns or lane changes and Longitudinal actions capture changes in speed, such as acceleration or deceleration. The separation of lateral and longitudinal actions in SDL reflects the distinct control systems in autonomous vehicles, allowing flexible and accurate representation of combined driving behaviors. This structure also aligns with established practices in vehicle dynamics, and allows us to evaluate autonomous vehicle responses across scenarios in a way that is interpretable, consistent, and comparable.
3.2. Hierarchical Levels of SDL Labels
@@ 4. Problem Statement
  2. (2)
    While real-world traffic involves agent interactions, this work isolates ego trajectories to demonstrate the feasibility of encoding vehicle behavior using SDLs. Future work can extend this to multi-agent scenarios (causal_agents, ).
  3. (3)
    We use the hierarchical SDL schema described in Section 3, though the method can adapt to frameworks like Scenic and OpenScenario.
  4. (4)
    We assume no ground truth action labels are provided for the trajectory. Without labeled data, supervised approaches like neural networks are infeasible. Unsupervised methods (e.g., k-means) rely on metrics such as ADE, which inadequately capture spatio-temporal behavior (Section 5).
@@
that defines a mapping between an input vehicle trajectory $\tau^{j}$, and produces an SDL label $L^{i}_{j}$ as the output describing the vehicle’s behavior. The SDL label $L^{i}_{j}=\langle\{L_{\text{lat},1},L_{\text{lat},2},\dots\},\{L_{\text{long},1},L_{\text{long},2},\dots\}\rangle$ is a tuple containing a chronologically ordered sequence of the lateral and longitudinal SDL labels as defined in Section 3.2. $f_{\Theta}$ is realized through a set of rules, where $\Theta$ denotes a set of separation thresholds. The problem of automatic SDL extraction is the problem of determining values for these thresholds.
4.1. Threshold Calculation Problem

### v2

@@ I Introduction
Autonomous vehicles (AVs) are nearing widespread deployment, but ensuring their safety remains a critical challenge. Public AV testing has revealed critical failures, sometimes leading to tragic outcomes [1, 2, 3, 4, 5, 6, 7, 8]. A methodical approach for measuring and improving safety for AVs is necessary to reduce these outcomes. A central question in AV development and testing is: How do we know which autonomous vehicle (AV) is safer? This fundamental question underlines much of the debate surrounding AV deployment and safety. Conventional metrics such as miles driven or disengagement counts provide a limited and often inadequate view of AV performance. Lower disengagements, for instance, does not necessarily equate to safer or more reliable driving behavior, as it fails to capture the full range of complex interactions AVs encounter in real-world traffic situations. To objectively compare the safety of two AVs, we need a framework that goes beyond these basic metrics.
@@
Comparing the safety of two AVs is challenging because they often operate under different conditions, different cities etc. Direct safety comparisons are unreliable without a common basis. Rather, a more viable approach is to analyze AV behavior when exposed to similar traffic scenarios. By evaluating AV performance in high-stakes situations, we gain deeper insight into their safety and robustness. However, implementing and scaling this approach requires the ability to automatically describe, interpret, and compare traffic scenarios across large datasets. This challenge motivates our work: How can we generate standardized, interpretable descriptions of traffic scenarios to identify both similar and rare situations for AV comparison?
@@
Scenario Description Languages (SDLs) provide structured, interpretable descriptions of traffic scenarios by capturing both static (road geometry, traffic signals, etc.) and dynamic (vehicle maneuvers) elements. SDLs can facilitate critical AV testing and safety tasks such as large-scale scenario discovery, real-to-sim transfers, post-deployment monitoring, and programmatic scenario generation. Additionally, by encoding traffic scenarios into SDL embeddings, we can compare and identify similar scenarios allowing us to observe AVs in a similar traffic situation and, hence, reasoning about their relative safety. Frameworks such as Scenic [9] and OpenScenario [10], offer ways to define SDLs for these purposes.
@@
However, despite their utility, significant limitations of constructing SDLs remain: Due to the lack of ground-truth annotations, the extraction of semantic SDL labels from AV data is currently manual, requiring intensive oversight. Manual labeling of such data is impractical due to the vast scale of AV trajectory datasets. Since AV data sets do not provide labeled actions, using supervised learning techniques to extract SDLs is infeasible. This bottleneck limits the scalability of SDLs for comprehensive scenario-based testing across diverse AV datasets. While advances in computer vision have made static element extraction (such as road geometry) more feasible [11, 12], extracting dynamic elements - particularly vehicle behaviors - is far more complex. Accurately capturing spatio-temporal actions requires more than video-based similarity; as it involves understanding nuanced trajectory patterns.
@@
To overcome these limitations, we introduce the Trajectory-to-Action Pipeline (TAP), a method designed to automate the extraction of interpretable vehicle behavior SDL descriptions from unlabeled AV trajectory data. TAP uses cross-entropy optimization to automatically derive SDL labels based on the dataset’s distribution. This automated approach addresses the pressing need for scalable, interpretable scenario descriptions, enabling TAP to compare vehicle behaviors and identify unique vehicle behaviors without the need for manual input. The contributions of this paper are as follows:
  1. 1.
@@ II Related Work
Scenario Description Languages Scenario Description Languages (SDLs) serve as high-level embeddings to describe traffic scenarios, capturing key elements like vehicles, pedestrians, and traffic lights. Existing SDLs such as Fortellix’s MSDL [13], Scenic [9], and OpenScenario [10] focus on comprehensive scene descriptions of environment, road geometry etc. However, our work diverges by concentrating specifically on describing autonomous vehicle motion to facilitate behavior comparison. Previous methods, such as Scenario2Vector [14, 15], have explored using SDLs for scenario similarity using video and natural language data. In contrast, our approach differs by introducing an automated pipeline for extracting vehicle motion - focused SDL labels, enhancing scalability and precision. Alternative approaches include graph-based representations like Roadscene2vec and GraphAD [16, 17, 18], which model scenes using nodes for agents and edges for their interactions. While effective for capturing vehicle interactions, these methods struggle to accurately represent vehicle motion dynamics. Recently, natural language descriptions [19, 20] have been used to capture complex scene interactions. Although expressive, this approach requires extensive training data and includes extraneous linguistic elements, making it less efficient for scenario comparison. Another approach, used by works such as iMotion [21] and MotionLM [22], uses Large Language Models to describe scenes. However, iMotion creates their prompts using template sentences, making it just another SDL. MotionLM treats motion tokens as a language, which loses the interpretability of a traditional SDL. In contrast, our method achieves a high information density while maintaining interpretability, focusing on the essential motion characteristics relevant to vehicle behavior.
@@
Scene Understanding Scene understanding is a vast field within computer vision, of which traffic scene understanding is a small portion. In [23], researchers use dashcam video for scenario retrieval. Other scene understanding work [24, 25, 26] focuses on modeling the relationship between road agents and their environment features like lanes, etc. For instance, Vid2Bev [26] converts camera footage into a bird’s-eye view, enabling AVs to create a vectorized world for planning and control. These algorithms are important for AVs to create a vectorized view of the world in which their planning and control modules operate. Since AV planning and control modules already operate in this vectorized space, we leverage the existing data available to the vehicle. This allows us to analyze vehicle behavior as perceived by the vehicle itself, rather than relying on camera-based perception. Thus, our focus is on analyzing vectorized vehicle data, bypassing the need to re-implement perception algorithms.
@@
Trajectory Similarity In the vectorized world used by AVs, vehicle behavior is represented by trajectories of positional history. Traditional methods for trajectory similarity, such as Average Displacement Error (ADE) [27] and Dynamic Time Warping (DTW) [28], compare distances between corresponding points. However, these methods struggle with varying trajectory lengths, requiring resampling, which can strip away crucial information. Deep learning methods [29] are more robust to length differences but remain sensitive to early deviations, leading to accumulated errors. In contrast, our Trajectory-to-Action Pipeline (TAP) uses Scenario Description Language (SDL) labels to capture interpretable action-level behaviors, preserving essential motion characteristics and ensuring robustness to trajectory variations.
III Scenario Description Language
@@
III-A Overview of SDL Structure
Figure 1: Actions in our SDL, organized along a spectrum of complexity. At the abstract level, there are three lateral and three longitudinal labels, which can be further subdivided into finer-grained labels for more detailed analysis.
SDLs provide a systematic way to describe driving behaviors. Based on the SDL framework outlined by [14], we define an SDL that characterizes autonomous vehicle behaviors through a defined set of temporal actions. Figure 1 illustrates the set of vehicle behaviors employed in our SDL framework. We partition the driving behavior of the autonomous vehicle in terms of lateral and longitudinal actions, where lateral actions capture directional movements, such as turns or lane changes and longitudinal actions capture changes in speed, such as acceleration or deceleration. The separation of lateral and longitudinal actions in SDL reflects the control systems in autonomous vehicles, allowing flexible and accurate representation of combined driving behaviors. This structure also aligns with established practices in vehicle dynamics, and allows us to evaluate autonomous vehicle responses across scenarios in a way that is interpretable, consistent, and comparable.
III-B Hierarchical Levels of SDL Labels
@@ IV Problem Statement
  2. 2.
    While real-world traffic involves agent interactions, this work isolates ego trajectories to demonstrate the feasibility of encoding vehicle behavior using SDLs. Future work can extend this to multi-agent scenarios [30].
  3. 3.
    We use the hierarchical SDL schema described in Section III, though the method can adapt to frameworks like Scenic and OpenScenario.
  4. 4.
    We assume no ground truth action labels are provided for the trajectory. Without labeled data, supervised approaches like neural networks are infeasible. Unsupervised methods (e.g., k-means) rely on measures such as ADE, which inadequately capture spatio-temporal behavior (Section VI-C).
@@
that defines a mapping between an input vehicle trajectory $\tau^{j}$, and produces an SDL label $L^{i}_{j}$ as the output describing the vehicle’s behavior. The SDL label $L^{i}_{j}=\langle\{L_{\text{lat},1},L_{\text{lat},2},\dots\},\{L_{\text{long},1},L_{\text{long},2},\dots\}\rangle$ is a tuple containing a chronologically ordered sequence of the lateral and longitudinal SDL labels as defined in Section III-B. $f_{\Theta}$ is realized through a set of rules, where $\Theta$ denotes a set of separation thresholds. The problem of automatic SDL extraction is the problem of determining values for these thresholds.
IV-A Threshold Calculation Problem